OpenStudy (anonymous):

lim as x approaches 0 -cos x + 1 - sin x all over x

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{ -\cos x + 1 - \sin x }{ x }\]

OpenStudy (anonymous):

apply l hopital rule5

OpenStudy (anonymous):